# Decentralized Autoregressive Generation

### v1

@@ 1 Introduction
The increase in performance of Multimodal Large Language Models (MLLMs) requires a massive increase in parameter counts (kaplan2020scalinglawsneurallanguage), which, in turn, requires enhanced infrastructure: training requires centralized clusters of thousands of GPUs with extensive gradient synchronization.
These developments present two critical challenges. First, training LLMs is inaccessible to academic institutions and small startups due to limited infrastructure. Second, the system is susceptible to individual node failures, which can disrupt the entire process (grattafiori2024llama3herdmodels). Existing mitigation strategies include aggressive checkpointing and related techniques.
A potential solution is a decentralized training approach, where multiple expert models are trained independently on disjoint subsets of the dataset, with only a few activated during inference time. This significantly reduces communication between nodes and enables academic institutions to collaborate. Recent studies (li2022branchtrainmergeembarrassinglyparalleltraining; gururangan2023scalingexpertlanguagemodels; no-need-to-talk) empirically demonstrate the efficiency of this approach for unimodal language modeling training in specific settings; however, existing work lacks a strict theoretical guarantee that decentralized training is equivalent to the centralized paradigm, making decentralized training a black box. Moreover, decentralized training of multimodal large language models has not yet been empirically explored.
mcallister2025decentralizeddiffusionmodels defined the Decentralized Flow Matching objective for Continuous Flow Matching (lipman2023flow) and demonstrated the equivalence between centralized and decentralized training of flow-matching models for image generation. At the same time, gat2024discrete introduced the Discrete Flow Matching theoretical framework, extending Continuous Flow Matching (lipman2023flow) to a discrete data domain. Inspired by these advancements, we derive the theoretical analysis of decentralized autoregressive generation.
@@
  • We extend the Discrete Flow Matching paradigm (gat2024discrete) to the discrete time domain, making the necessary modifications and defining the corresponding conditions.
  • Similar to mcallister2025decentralizeddiffusionmodels, we define the Decentralized Discrete Flow Matching objective and show that the probability-generating velocity is a linear combination of expert flows.
@@
  • We conduct experiments demonstrating the equivalence between decentralized and centralized training for multimodal language models in two distinct settings: LLaVA (liu2024improvedbaselinesvisualinstruction) and InternVL (chen2025expandingperformanceboundariesopensource).
@@ 2.1 Continuous Time Discrete Flow Matching
Discrete Flow Matching (gat2024discrete) is a theoretical framework that describes discrete generative modeling processes in terms similar to Continuous Flow Matching (lipman2023flow). Namely, having vocabulary or tokens $[d]$, source distribution $X_{0}$ with PMF $p(x)$ of token sequences $x=(x^{1},\ldots,x^{N})\in[d]^{N}$ and target distribution $X_{1}$ with PMF $q(x)$, we define the probability path $p_{t}(x),t\in[0,1]$ s.t. $p_{0}(x)=p(x)$, $p_{1}(x)=q(x)$ as follows:
@@
As shown in (gat2024discrete), if the continuity equation holds, then the probability generating velocity generates a probability path as defined in (7-8).
@@ 3.2 Autoregressive sampling
To define autoregressive sampling, we define the source and target distribution as a special case of C-coupling from (gat2024discrete):
@@ 3.3 Decentralization
The above formalization automatically makes Theorem 1 true for autoregressive sampling. Now, similar to (mcallister2025decentralizeddiffusionmodels), if we define a target distribution $X_{1}$ as a union of disjoint clusters $S_{k},\quad k\in\{1,2,\dots,K\}$, we can rewrite Equation 24 as follows:
@@ 4 Method
Since the effectiveness of decentralized discrete flow matching in case of pure LM has already been shown in previous work (no-need-to-talk; gururangan2023scalingexpertlanguagemodels), we focus on Vision Language Models as another instance of autoregressive sampling.
@@ 5.1 LLaVA
In the first part of the experiments, we utilize the LLaVA-1.5 (liu2024improvedbaselinesvisualinstruction) fine-tuning dataset, consisting of 625k image-text samples and 41k text-only samples from diverse sources.
@@ 5.2 InternVL
In the second phase of our experiments, we utilize a subset of the InternVL 2.5 (chen2025expandingperformanceboundariesopensource) Stage 2 fine-tuning data mixture (see Table 1). We curated datasets representing diverse task types to ensure comprehensive coverage.
@@
Captioning TextCaps (sidorov2020textcapsdatasetimagecaptioning) , ShareGPT4o (cui2025comprehensive)
General QA VQAv2(balanced_vqa_v2), GQA(hudson2019gqanewdatasetrealworld)
Chart ChartQA(masry2022chartqabenchmarkquestionanswering)
OCR InfoVQA(mathew2021infographicvqa) , TextVQA(singh2019vqamodelsread)
Knowledge KVQA(shahMYP19), A-OKVQA(schwenk2022aokvqabenchmarkvisualquestion)
Document DocVQA(mathew2021docvqadatasetvqadocument)
Grounding RefCOCO/+/g (mao2016generationcomprehensionunambiguousobject; yu2016modelingcontextreferringexpressions)
Science AI2DKembhavi2016ADI, ScienceQAlu2022learn
Conversation ALLaVA(chen2024allavaharnessinggpt4vsynthesizeddata)
@@ 5.3 Evaluation
We evaluate our LLaVA models across total of 8 benchmarks, used in the liu2024improvedbaselinesvisualinstruction. For InternVL, we evaluate across 14 benchmarks, aiming to assess the models in the general question answering, OCR and other tasks, used in the chen2025expandingperformanceboundariesopensource.
@@ 5.5 Ablation Study
In the main InternVL experiments, we use spherical balanced k-means clustering algorithm for dataset partitioning and routing. To test the impact of the clustering algorithm, we also train $K=2$ experts using 2-stage balanced spherical k-means algorithm inspired by mcallister2025decentralizeddiffusionmodels. Namely, first stage is fine clustering into $k=1024$ clusters in unbalanced manner. Second stage is coarse clustering of fine centroids into $K=2$ coarse clusters in a balanced manner. As shown in Table 9, 2-stage k-means leads to marginally worse general QA performance (-0.37 VQAv2), at the same time improves OCR performance (+0.78 TextVQA).

### v2

@@ 1 Introduction
The increase in performance of Multimodal Large Language Models (MLLMs) requires a massive increase in parameter counts (Kaplan et al., 2020), which, in turn, requires enhanced infrastructure: training requires centralized clusters of thousands of GPUs with extensive gradient synchronization.
These developments present two critical challenges. First, training LLMs is inaccessible to academic institutions and small startups due to limited infrastructure. Second, the system is susceptible to individual node failures, which can disrupt the entire process (Grattafiori et al., 2024). Existing mitigation strategies include aggressive checkpointing and related techniques.
A potential solution is a decentralized training approach, where multiple expert models are trained independently on disjoint subsets of the dataset, with only a few activated during inference time. This significantly reduces communication between nodes and enables academic institutions to collaborate. Recent studies (Li et al., 2022; Gururangan et al., 2023; Filippova et al., 2025) empirically demonstrate the efficiency of this approach for unimodal language modeling training in specific settings; however, existing work lacks a strict theoretical guarantee that decentralized training is equivalent to the centralized paradigm, making decentralized training a black box. Moreover, decentralized training of multimodal large language models has not yet been empirically explored.
McAllister et al. (2025) defined the Decentralized Flow Matching objective for Continuous Flow Matching (Lipman et al., 2023) and demonstrated the equivalence between centralized and decentralized training of flow-matching models for image generation. At the same time, Gat et al. (2024) introduced the Discrete Flow Matching theoretical framework, extending Continuous Flow Matching (Lipman et al., 2023) to a discrete data domain. Inspired by these advancements, we derive the theoretical analysis of decentralized autoregressive generation.
@@
  • We extend the Discrete Flow Matching paradigm (Gat et al., 2024) to the discrete time domain, making the necessary modifications and defining the corresponding conditions.
  • Similar to McAllister et al. (2025), we define the Decentralized Discrete Flow Matching objective and show that the probability-generating velocity is a linear combination of expert flows.
@@
  • We conduct experiments demonstrating the equivalence between decentralized and centralized training for multimodal language models in two distinct settings: LLaVA (Liu et al., 2024) and InternVL (Chen et al., 2025).
@@ 2.1 Continuous Time Discrete Flow Matching
Discrete Flow Matching (Gat et al., 2024) is a theoretical framework that describes discrete generative modeling processes in terms similar to Continuous Flow Matching (Lipman et al., 2023). Namely, having vocabulary or tokens $[d]$, source distribution $X_{0}$ with PMF $p(x)$ of token sequences $x=(x^{1},\ldots,x^{N})\in[d]^{N}$ and target distribution $X_{1}$ with PMF $q(x)$, we define the probability path $p_{t}(x),t\in[0,1]$ s.t. $p_{0}(x)=p(x)$, $p_{1}(x)=q(x)$ as follows:
@@
As shown in (Gat et al., 2024), if the continuity equation holds, then the probability generating velocity generates a probability path as defined in (7-8).
@@ 3.2 Autoregressive sampling
To define autoregressive sampling, we define the source and target distribution as a special case of C-coupling from (Gat et al., 2024):
@@ 3.3 Decentralization
The above formalization automatically makes Theorem 1 true for autoregressive sampling. Now, similar to (McAllister et al., 2025), if we define a target distribution $X_{1}$ as a union of disjoint clusters $S_{k},\quad k\in\{1,2,\dots,K\}$, we can rewrite Equation 24 as follows:
@@ 4 Method
Since the effectiveness of decentralized discrete flow matching in case of pure LM has already been shown in previous work (Filippova et al., 2025; Gururangan et al., 2023), we focus on Vision Language Models as another instance of autoregressive sampling.
@@ 5.1 LLaVA
In the first part of the experiments, we utilize the LLaVA-1.5 (Liu et al., 2024) fine-tuning dataset, consisting of 625k image-text samples and 41k text-only samples from diverse sources.
@@ 5.2 InternVL
In the second phase of our experiments, we utilize a subset of the InternVL 2.5 (Chen et al., 2025) Stage 2 fine-tuning data mixture (see Table 1). We curated datasets representing diverse task types to ensure comprehensive coverage.
@@
Captioning TextCaps (Sidorov et al., 2020) , ShareGPT4o (Cui et al., 2024)
General QA VQAv2(Goyal et al., 2017), GQA(Hudson and Manning, 2019)
Chart ChartQA(Masry et al., 2022)
OCR InfoVQA(Mathew et al., 2021a) , TextVQA(Singh et al., 2019)
Knowledge KVQA(Sanket Shah and Talukdar, 2019), A-OKVQA(Schwenk et al., 2022)
Document DocVQA(Mathew et al., 2021b)
Grounding RefCOCO/+/g (Mao et al., 2016; Yu et al., 2016)
Science AI2DKembhavi et al. (2016), ScienceQALu et al. (2022)
Conversation ALLaVA(Chen et al., 2024)
@@ 5.3 Evaluation
We evaluate our LLaVA models across total of 8 benchmarks, used in the Liu et al. (2024). For InternVL, we evaluate across 14 benchmarks, aiming to assess the models in the general question answering, OCR and other tasks, used in the Chen et al. (2025).
@@ 5.5 Ablation Study
In the main InternVL experiments, we use spherical balanced k-means clustering algorithm for dataset partitioning and routing. To test the impact of the clustering algorithm, we also train $K=2$ experts using 2-stage balanced spherical k-means algorithm inspired by McAllister et al. (2025). Namely, first stage is fine clustering into $k=1024$ clusters in unbalanced manner. Second stage is coarse clustering of fine centroids into $K=2$ coarse clusters in a balanced manner. As shown in Table 9, 2-stage k-means leads to marginally worse general QA performance (-0.37 VQAv2), at the same time improves OCR performance (+0.78 TextVQA).